# Augmented Lagrangian Solvers for Poroelasticity with Fracture Contact Mechanics

### v1

@@ 2.4 Boundary conditions and summary
On external boundaries, we impose either Dirichlet or Neumann boundary conditions. The mechanical boundary data is accordingly either prescribed values of displacement $\bm{u}$ or traction $\sigma\cdot\bm{n}$, and fluid boundary data is either prescribed values of pressure $p$ or mass flux $\rho\bm{v}\cdot\bm{n}$. Finally, we impose zero fluid mass flux on immersed fracture tips, supported by the discussion in [33].
@@ 3 Augmented Lagrangian solvers
In this section, we will present three different nonlinear solvers for the system of equations presented in the previous section. All three solvers are based on the augmented Lagrangian formulation of the frictional contact problem. Two of these solvers are classical, namely the generalized Newton method and the implicit return map, while the third solver is new. To ease the presentation and stress their agnostic character to spatial discretizations, the solvers will be presented in a semi-discrete setting, where the equations are assumed to be discretized in time, using the implicit Euler method, but continuous in space. As a consequence, the slip velocity $\dot{\bm{u}}]]_{\tau}$ now refers to the time increment of the tangential displacement jump, i.e. $\dot{\bm{u}}]]_{\tau}=[[\bm{u}]]_{\tau}-[[\bm{u}]]_{\tau}^{\text{prev}}$, where ”prev” denotes the previous time step, and the time step size is implicitly subsumed in the scaling constant in (17). We will give some details in section 3.4 of the particular spatial discretizations used in the numerical experiments of the present work.
Notation. We denote nonlinear iteration indices by superscripts. Moreover, it will be convenient to separate the contact conditions (15), (17) from the rest of the equation system, as the contact conditions are treated differently, depending on the solver, while the residual system remains unchanged. We denote the residual system by $G$, and collect all unknowns into a vector $\bm{x}$. Hence, the equation
@@
describes the semi-discrete formulations of all equations from table 1, with the exception of (15) and (17), at a particular time step.
@@ 3.1 The generalized Newton method
The nonpenetration condition (15) and Coulomb friction law (17) can be reformulated as the zero sets of certain nonsmooth functions called complementarity functions. Such a reformulation can be derived from a theoretical consideration of the augmented Lagrangian formulation of the contact problem, as shown in [25]. There are several ways of formulating these complementarity functions [25, 26, 27]. In the present work, we employ the formulation of Hüeber et al. [27], in which it is shown that, for every augmentation parameter $c>0$, the nonpenetration condition (15) is equivalent to:
@@
Here $b=-F\lambda_{n}$ is the friction bound and $\chi$ is a characteristic function that equals 1 if $b\leq 0$ and 0 if $b>0$, thus enforcing $\bm{\lambda}_{\tau}=\bm{0}$ in the case of no contact. By replacing (15) and (17) with (23) and (24), the entire equation system of table 1 may be written compactly in residual form. Using the previously defined residual system (22), the full system of equations to be solved at a specific time step reads
@@
This system is only differentiable almost everywhere (in the sense of the Lebesgue measure), due to the non-smoothness of the complementarity functions. However, the non-differentiable points can easily be handled by employing generalized derivatives, using the definition found in Clarke [34]. A set of generalized Jacobians for the system (25) at a point $\bm{x}^{*}$ is defined as follows:
@@
where $\Omega_{F}$ is the set of points on which $F$ is differentiable, with corresponding Jacobian $DF$, and ”conv” denotes the convex hull. Using these generalized Jacobians, a corresponding generalized Newton method can be defined [25, 27], which is described below. We will abbreviate the method by GNM.
@@
1. Pick an initial guess $\bm{x}^{(0)}$. Set $k=0$.
@@ 3. Convergence check:
  if $\bm{x}^{(k+1)}$ fulfills convergence criterion then
@@
Let us briefly comment on the augmentation parameter $c$ appearing in the complementarity functions. Since the two complementarity functions are an exact reformulation of the nonpenetration condition (15) and Coulomb friction law (17) for all $c>0$, this parameter may be chosen freely without affecting the accuracy of the solution. This reflects the fact that the solution to the augmented Lagrangian formulation is exact for all positive values of the augmentation parameter, as mentioned in the introduction. It is important to note, however, that the choice of the parameter can greatly affect the performance of the generalized Newton method (this is also the case for the other nonlinear solvers considered in this work); this will be shown in the numerical experiments of section 4.
Following [27, 7], we define ”generalized” contact states, based on the arguments of the maximum functions of (23) and (24). We group open, stick and slip states as follows:
Open $=\{-\lambda_{n}-c([[\bm{u}]]_{n}-g)\leq 0\},$ (28)
Stick $=\{b>0\ \text{and}\ \lVert\bm{\lambda}_{\tau}+c[[\dot{\bm{u}}]]_{\tau}\rVert<b\},$
Slip $=\{b>0\ \text{and}\ \lVert\bm{\lambda}_{\tau}+c[[\dot{\bm{u}}]]_{\tau}\rVert\geq b\}.$
The physical contact states of open, stick and slip, as described by the nonpenetration condition (15) and Coulomb friction law (17), are subsets of the corresponding sets above, hence the notion of generalized contact states. References to the sets (28) will be made in section 4.
@@ 3.2 The implicit return map method
Return map methods are commonly used in computational plasticity [35, 36], and have also be applied to frictional contact problems [37, 38, 28]. The defining feature of such methods is that for every iteration, a trial value is proposed for a variable subject to constraints, and violation of these constraints is checked for this value. If the trial value lies outside the feasible set, it is projected back to the boundary of the set, this projection is what is called the return map. This methodology was first applied to the augmented Lagrangian formulation of frictional contact mechanics in [28].
@@
$\lambda_{n}^{\text{trial}}$ $=\lambda_{n}^{(k)}+c([[\bm{u}^{(k+1)}]]_{n}-g(\bm{u}^{(k+1)})),$ (29a)
$\bm{\lambda}_{\tau}^{\text{trial}}$ $=\bm{\lambda}_{\tau}^{(k)}+c[[\dot{\bm{u}}^{(k+1)}]]_{\tau},$ (29b)
for some $c>0$. These formulas result in the contact traction converging in the case when $\bm{u}]]_{n}-g\to 0$ and $\dot{\bm{u}}]]_{\tau}\to\bm{0}$ (stick), while the remaining cases (open, slip) are handled by the return map. Some new notation is introduced at this point: We separate the contact traction $\bm{\lambda}$ from the rest of the unknowns, and denote the ”residual unknowns” by $\bm{r}$. As we have previously defined the total collection of unknowns as $\bm{x}$, we may write $\bm{x}=(\bm{r},\bm{\lambda)}$. The residual system (22) is accordingly rewritten as $G(\bm{r,\bm{\lambda)}}=\bm{0}$. Using this notation, we formulate the return map method below, which is a variant of the method from [28]. We emphasize that the return map is only performed implicitly in the method (this point will be expanded upon), hence we call it the implicit return map method (IRM).
@@
1. Pick an initial guess $\bm{x}^{0}=(\bm{r}^{0},\bm{\lambda}^{0})$. Set $k=0.$
2. Solve the following system for $(\bm{r}^{(k+1)},\bm{\lambda}^{(k+1)}):$
@@
  where $b^{(k+1)}=-F\lambda_{n}^{(k+1)}$.
@@ 3. Convergence check:
  if $(\bm{r}^{(k+1)},\bm{\lambda}^{(k+1)})$ fulfills convergence criterion then
@@
Note that algorithm 2 defines a nested loop; each outer iteration $k$ corresponds to an inner loop for solving the (in general nonlinear) system (30) together with the return map (31). We observe that the return map is implicit due to the presence of $\bm{\lambda}^{(k+1)}$ in (30), meaning that (30) and (31) must be solved simultaneously.
@@
$\lambda_{n}^{(k+1)}+\max(0,-\lambda_{n}^{(k)}-c([[\bm{u}^{(k+1)}]]_{n}-g(\bm{u}^{(k+1)})))=0,$ (33a)
$\chi^{(k+1)}\bm{\lambda}_{\tau}^{(k+1)}+(1-\chi^{(k+1)})\left[b^{(k+1)}(\bm{\lambda}_{\tau}^{(k)}+c[[\dot{\bm{u}}^{(k+1)}]]_{\tau})-\max(b^{(k+1)},\lVert\bm{\lambda}_{\tau}^{(k)}+c[[\dot{\bm{u}}^{(k+1)}]]_{\tau}\rVert)\bm{\lambda}_{\tau}^{(k+1)}\right]=\bm{0}.$ (33b)
Equations (33a) and (33b) are nearly identical to the complementarity functions (23) and (24); the only difference is the terms involving the traction at the previous iteration, $\bm{\lambda}^{(k)}$, which is a known quantity in these equations. One can interpret (33a) and (33b) as regularized contact conditions, in the sense that they convert the contact conditions from graphs to functions, due to the $\bm{\lambda}^{(k)}$-terms. Hence, IRM is equivalent to solving a sequence of systems with regularized contact conditions. More precisely, for a fixed $\bm{\lambda}^{(k)}$ one may define the regularized complementarity functions
$\mathcal{C}_{n}^{(k)}$ $:=\lambda_{n}+\max(0,-\lambda_{n}^{(k)}-c([[\bm{u}]]_{n}-g(\bm{u}))),$
$\mathcal{C}_{\tau}^{(k)}$ $:=\chi\bm{\lambda}_{\tau}+(1-\chi)\left[b(\bm{\lambda}_{\tau}^{(k)}+c[[\dot{\bm{u}}]]_{\tau})-\max(b,\lVert\bm{\lambda}_{\tau}^{(k)}+c[[\dot{\bm{u}}]]_{\tau}\rVert)\bm{\lambda}_{\tau}\right].$
@@
These regularized systems may be solved by GNM (algorithm 1), and this is how IRM has been implemented in practice in the numerical experiments of section 4. We note also that the return map (31) is equivalent to updating the contact tractions (or augmented Lagrange multipliers) by fixed-point iteration, according to (32). This corresponds to the classical Uzawa algorithm [22]. Hence, IRM can also be interpreted as an Uzawa algorithm. Again we emphasize that the fixed-point iteration is only performed implicitly in IRM, resulting in an implicit Uzawa algorithm. It is known that for purely mechanical contact problems, implicit Uzawa algorithms converge for a larger range of augmentation parameters than their explicit counterparts [29], at the cost of needing to solve a larger and more complicated system.
The augmentation parameter $c$ takes on a different meaning for IRM, compared to GNM. Here it directly controls the strength of the regularization, and as $c\to\infty$, the regularized complementarity functions (33a) and (33b) converge to the exact functions (23) and (24). However, like GNM, the augmented Lagrangian formulation ensures that IRM also converges to the non-regularized solution for any $c>0$, although in this case the convergence speed is highly dependent on the value of $c$, with lower values requiring more iterations of the outer loop to converge. This is a well-known feature of Uzawa algorithms [24, 29].
Similarly to how the sets (28) describing the different contact states could be defined in relation to the original complementarity functions, we can for each outer iteration $k$ of IRM define the ”regularized” contact states,
Regularized open $=\{-\lambda_{n}^{(k)}-c([[\bm{u}^{(k+1)}]]_{n}-g(\bm{u}^{(k+1)})\leq 0\},$ (35)
Regularized stick $=\{b^{(k+1)}>0\ \text{and}\ \lVert\bm{\lambda}_{\tau}^{(k)}+c[[\dot{\bm{u}}^{(k+1)}]]_{\tau}\rVert<b^{(k+1)}\},$
Regularized slip $=\{b^{(k+1)}>0\ \text{and}\ \lVert\bm{\lambda}_{\tau}^{(k)}+c[[\dot{\bm{u}}^{(k+1)}]]_{\tau}\rVert\geq b^{(k+1)}\},$
corresponding to the different arguments of the maximum functions of the regularized complementarity functions (33a) and (33b). References to these sets will be made in the numerical experiments of section 4.
@@ 3.3 A generalized Newton method with a return map
In the previous section, we showed a connection between the implicit return map method, the generalized Newton method and the complementarity functions. These observations provide the motivation for a new type of nonlinear solver, which we propose in this section. The idea is to design a quasi-Newton method that incorporates the return map (31) of the contact tractions, but without needing to regularize the contact conditions. This can be achieved if, rather than defining the return map implicitly through regularization of the complementarity functions, as done in IRM, we keep the original complementarity functions, and instead perform the return map (31) explicitly after applying an iteration of algorithm 1 (GNM). In place of (29), the trial traction is now defined as the newly obtained guess of the traction from applying one iteration of GNM. Performing the return map in this way removes the need to solve a sequence of regularized systems, and hence also removes the nested loop of IRM. The resulting nonlinear solver is described below, using the notation introduced in section 3.2. We call it the generalized Newton method with a return map, and abbreviate it by GNM-RM.
@@
1. Pick an initial guess $\bm{x}^{0}=(\bm{r}^{0},\bm{\lambda}^{0})$. Set $k=0$.
2. Using the generalized Newton method, perform one iteration on the system (25) to get
   $(\bm{r}^{(k+1)},\bm{\lambda}^{\text{trial}}).$

### v2

@@ 2.4 Boundary conditions and summary
On external boundaries, we impose either Dirichlet or Neumann boundary conditions. The mechanical boundary data is accordingly either prescribed values of displacement $\bm{u}$ or traction $\sigma\cdot\bm{n}$, and fluid boundary data is either prescribed values of pressure $p$ or mass flux $\rho\bm{v}\cdot\bm{n}$. Finally, we impose zero fluid mass flux on immersed fracture tips, supported by the discussion in [both2024].
@@ 3 Augmented Lagrangian solvers
In this section, we will present three different nonlinear solvers for the system of equations presented in the previous section. All three solvers are based on the augmented Lagrangian formulation of the frictional contact problem. Two of these solvers are classical, namely the generalized Newton method and the implicit return map, while the third solver is new. To ease the presentation and stress their agnostic character to spatial discretizations, the solvers will be presented in a semi-discrete setting, where the equations are assumed to be discretized in time, using the implicit Euler method, but continuous in space. We will give some details in section 3.4 of the particular spatial discretizations used in the numerical experiments of the present work.
Notation. The full set of primary unknowns is collected into a vector $\bm{x}$, consisting of the pressure, displacement, interface Darcy flux, interface displacement and contact traction, as outlined in table 1. Nonlinear iteration indices will be denoted by a superscript, i.e. $\bm{x}^{(k)}$. Moreover, we will separate the contact conditions, (16) and (18), from the rest of the equation system. This is done because the implicit return map method enforces the contact conditions differently from the other two solvers, while the remaining equations are treated the same by all solvers. We denote by $G$ the residual system in which the contact conditions have been excluded. Hence, the equation
@@
constitutes the semi-discrete formulations of equations (2), (3), (8) and (14), or in other words, every equation of table 1 with the exception of the nonpenetration conditon (16) and Coulomb friction law (18).
@@ 3.1 The generalized Newton method
The nonpenetration condition (16) and Coulomb friction law (18) can be reformulated as the zero sets of certain nonsmooth functions called complementarity functions. Such a reformulation can be derived from a theoretical consideration of the augmented Lagrangian formulation of the contact problem, as shown in [ALART1991353]. There are several ways of formulating these complementarity functions [ALART1991353, klarbring, hueber]. In the present work, we employ the formulation of Hüeber et al. [hueber], in which it is shown that, for every augmentation parameter $c>0$, the nonpenetration condition (16) is equivalent to:
@@
Here $b=-\mu\lambda_{n}$ is the friction bound and $\chi$ is a characteristic function that equals 1 if $\lvert b\rvert\leq 0$ and 0 if $\lvert b\rvert>0$, thus enforcing $\bm{\lambda}_{\tau}=\bm{0}$ in the case of no contact. Note that, as the equations in this section are assumed to be discretized in time, the slip velocity $\dot{\bm{u}}]]_{\tau}$ now refers to the time increment of the tangential displacement jump. In other words, we have $\dot{\bm{u}}]]_{\tau}=[[\bm{u}]]_{\tau}-[[\bm{u}]]_{\tau}^{\text{prev}}$, where “prev” denotes the previous time step, and the time step size is implicitly subsumed in the augmentation parameter $c$.
By replacing (16) and (18) with (26) and (27), the entire equation system of table 1 may be written compactly in residual form. Using the previously defined residual system (25), the full system of equations to be solved at a specific time step reads
@@
This system is only differentiable almost everywhere (in the sense of the Lebesgue measure), due to the non-smoothness of the complementarity functions. However, the non-differentiable points can easily be handled by employing generalized derivatives, using the definition found in Clarke [clarke]. A set of generalized Jacobians for the system (28) at a point $\bm{x}^{*}$ is defined as follows:
@@
where $\Omega_{F}$ is the set of points on which $F$ is differentiable, with corresponding Jacobian $DF$, and “conv” denotes the convex hull. Using these generalized Jacobians, a corresponding generalized Newton method can be defined, with the resulting algorithm being equivalent to a primal-dual active set method [hintermuller2002primal, hueber]. We describe this generalized Newton method below, and abbreviate it by GNM.
@@
1. Pick an initial guess $\bm{x}^{(0)}$. Choose tolerances TOL1 and TOL2. Set $k=0$.
@@ 3. Convergence check:
  if $\lVert\bm{x}^{(k+1)}-\bm{x}^{(k)}\rVert<\text{TOL1}$ and $\lVert F(\bm{x}^{(k+1)}\rVert<\text{TOL2}$ then
@@
Let us briefly comment on the augmentation parameter $c$ appearing in the complementarity functions. Since the two complementarity functions are an exact reformulation of the nonpenetration condition (16) and Coulomb friction law (18) for all $c>0$, this parameter may in theory be chosen freely without affecting the accuracy of the solution. However, on a practical level, the choice of the parameter can greatly affect the performance of the generalized Newton method (this is also the case for the other nonlinear solvers considered in this work), and this will be an important point in the numerical experiments of section 4.
Note also that the augmentation parameter can be chosen differently for the normal (26) and tangential (27) complementarity functions. However, we will for simplicity assume throughout this paper that it is equal for the two complementarity functions; this will also hold for the subsequent algorithms of sections 3.2 and 3.3. In our numerical experiments, we did not find any significant enough advantages in using two different parameters to warrant the added complexity and bookkeeping.
@@ 3.2 The implicit return map method
Return map methods are commonly used in computational plasticity [comp_plasticity, computational_inelasticity], and have also been applied to frictional contact problems [GIANNAKOPOULOS1989157, wriggers_return_map, SIMO199297]. The defining feature of such methods is that for every iteration, a trial value is proposed for a variable subject to constraints, and violation of these constraints is checked for this value. If the trial value lies outside the feasible set, it is projected back to the boundary of the set; this projection is what is called the return map. This methodology was first applied to the augmented Lagrangian formulation of frictional contact mechanics in [SIMO199297].
@@
$\lambda_{n}^{\text{trial}}$ $=\lambda_{n}^{(k)}+c^{(k)}([[\bm{u}^{(k+1)}]]_{n}-g(\bm{u}^{(k+1)})),$ (31a)
$\bm{\lambda}_{\tau}^{\text{trial}}$ $=\bm{\lambda}_{\tau}^{(k)}+c^{(k)}[[\dot{\bm{u}}^{(k+1)}]]_{\tau},$ (31b)
for some $c^{(k)}>0$, which may be chosen differently for each iteration. These formulas result in the contact traction converging in the case when $\bm{u}]]_{n}-g\to 0$ and $\dot{\bm{u}}]]_{\tau}\to\bm{0}$ (stick), while the remaining cases (open, slip) are handled by the return map. Some new notation is introduced at this point: We separate the contact traction $\bm{\lambda}$ from the rest of the unknowns, and denote the remaining unknowns by $\bm{r}$. As we have previously defined the total collection of unknowns as $\bm{x}$, we may write $\bm{x}=(\bm{r},\bm{\lambda})$. The residual system (25) is accordingly rewritten as $G(\bm{r,\bm{\lambda}})=\bm{0}$. Using this notation, we formulate the return map method below, which is equivalent to the method from [SIMO199297]. We refer to it as the implicit return map method (IRM).
@@
1. Pick an initial guess $\bm{x}^{0}=(\bm{r}^{0},\bm{\lambda}^{0})$. Choose tolerances TOL1 and TOL2. Set $k=0.$
2. Solve the following system for $\bm{x}^{(k+1)}=(\bm{r}^{(k+1)},\bm{\lambda}^{(k+1)}):$
@@
  where $b^{(k+1)}=-\mu\lambda_{n}^{(k+1)}$, and $\lambda_{n}^{\text{trial}}$, $\bm{\lambda}_{\tau}^{\text{trial}}$ are given by (31).
@@ 3. Convergence check:
  if $\lVert\bm{x}^{(k+1)}-\bm{x}^{(k)}\rVert<\text{TOL1}$ and $\lVert F(\bm{x}^{(k+1)})\rVert<\text{TOL2}$ then
@@
As alluded to by the naming of the algorithm, the return map is only performed implicitly, due to the presence of $\bm{\lambda}^{(k+1)}$ in (32). This means that (32) and (33) must be solved simultaneously. We note also that algorithm 2 defines a nested loop; each outer iteration $k$ consists of solving the system defined by (32) and (33). This is a nonlinear system, which must be solved in a separate inner loop. The outer loop can be interpreted as the augmentation procedure of the Lagrange multipliers (contact tractions), in addition to re-solving the residual system with the updated Lagrange multipliers. Note also that the convergence check is done on the system (28) defined earlier, thus ensuring that the contact constraints are adequately satisfied by checking the residuals of the complementarity functions (26) and (27).
@@
$\lambda_{n}^{(k+1)}+\max(0,-\lambda_{n}^{(k)}-c^{(k)}([[\bm{u}^{(k+1)}]]_{n}-g(\bm{u}^{(k+1)})))=0,$ (35a)
$\chi^{(k+1)}\bm{\lambda}_{\tau}^{(k+1)}+(1-\chi^{(k+1)})\left[b^{(k+1)}(\bm{\lambda}_{\tau}^{(k)}+c^{(k)}[[\dot{\bm{u}}^{(k+1)}]]_{\tau})-\max(b^{(k+1)},\lVert\bm{\lambda}_{\tau}^{(k)}+c^{(k)}[[\dot{\bm{u}}^{(k+1)}]]_{\tau}\rVert)\bm{\lambda}_{\tau}^{(k+1)}\right]=\bm{0}.$ (35b)
Equations (35a) and (35b) are nearly identical to the complementarity functions (26) and (27); the only differences are the terms involving the traction at the previous iteration, $\bm{\lambda}^{(k)}$, which is a known quantity in these equations. One can interpret (35a) and (35b) as regularized contact conditions, in the sense that they convert the contact conditions from graphs to functions, due to the $\bm{\lambda}^{(k)}$-terms. Hence, IRM is equivalent to solving a sequence of systems with regularized contact conditions. More precisely, for a fixed $\bm{\lambda}^{(k)}$ one may define the regularized complementarity functions
$\mathcal{C}_{n}^{(k)}$ $:=\lambda_{n}+\max(0,-\lambda_{n}^{(k)}-c^{(k)}([[\bm{u}]]_{n}-g(\bm{u}))),$
$\mathcal{C}_{\tau}^{(k)}$ $:=\chi\bm{\lambda}_{\tau}+(1-\chi)\left[b(\bm{\lambda}_{\tau}^{(k)}+c^{(k)}[[\dot{\bm{u}}]]_{\tau})-\max(b,\lVert\bm{\lambda}_{\tau}^{(k)}+c^{(k)}[[\dot{\bm{u}}]]_{\tau}\rVert)\bm{\lambda}_{\tau}\right].$
@@
These regularized systems may be solved by GNM (algorithm 1), and this is how IRM has been implemented in practice in the numerical experiments of section 4. We note also that the return map (33) is equivalent to updating the contact tractions (or augmented Lagrange multipliers) by fixed-point iteration, according to (34). This corresponds to the classical Uzawa algorithm [WriggersPeter2006CCM]. Hence, IRM can also be interpreted as an Uzawa algorithm. Again we emphasize that the fixed-point iteration is only performed implicitly in IRM, resulting in an implicit Uzawa algorithm. It is known that for purely mechanical contact problems, implicit Uzawa algorithms converge for a larger range of augmentation parameters than their explicit counterparts [Kunisch_Stadler_2005], at the cost of needing to solve a larger and more complicated system.
The augmentation parameter $c$ takes on a different meaning for IRM, compared to GNM. Here it directly controls the strength of the regularization, and as $c\to\infty$, the regularized complementarity functions (35a) and (35b) converge to the exact functions (26) and (27). However, the augmentation procedure on the contact tractions ensures that IRM will in theory converge to the non-regularized solution for any $c>0$, just like GNM. That being said, IRM still differs from GNM in that the convergence speed is highly dependent on the value of $c$, with lower values requiring more iterations of the outer loop to converge. This is a well-known feature of Uzawa algorithms [RenardSurvey, Kunisch_Stadler_2005]. Speedup of the outer loop may be achieved by gradually increasing $c$ with increasing outer loop iterations. Since $c$ need not be taken to infinity to obtain convergence, one may define an upper limit for its value and stop the increase once this limit is reached.
@@ 3.3 A generalized Newton method with a return map
In this section, we will present a new augmented Lagrangian solver, which combines features of both previous solvers. The central idea of this solver is to perform the same generalized Newton iterations as GNM, but after each iteration, a return map of the contact tractions is performed before proceeding to the next iteration. This is motivated by the fact that Newton solvers are well-known to often display divergent behavior (e.g. cycling) outside the radius of quadratic convergence. Performing a return map after each iteration might help to stabilize the solver in such cases. The solver is described below, using the notation introduced in section 3.2. We refer to it as the generalized Newton method with a return map, and abbreviate it by GNM-RM.
@@
1. Pick an initial guess $\bm{x}^{0}=(\bm{r}^{0},\bm{\lambda}^{0})$. Choose tolerances TOL1 and TOL2. Set $k=0$.
2. Perform one generalized Newton iteration (30) on the system (28) to get
  $\bm{x}^{\text{trial}}=(\bm{r}^{(k+1)},\bm{\lambda}^{\text{trial}}).$